### v2

@@ Appendix A Proofs
$\overline{\mu}_{i}^{I}$ $\coloneqq\mathbb{E}\left[\left.G_{V}\left(S_{j}\right)\tau^{{}^{\prime}}\left(V_{j}\right)p_{j}q_{j}^{{}^{\prime}}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\ u_{i}\coloneqq Y_{i}-G\left(S_{i}\right),$
@@
$\widehat{\Omega}_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{{}^{\prime}}\widehat{P}^{-1}\left(\widehat{\Sigma}+\widehat{\Sigma}_{1}\right)\widehat{P}^{-1}\overline{p}^{M_{2}}\left(v,w\right),$ (A.14)
$\widehat{\Sigma}$ $\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{p}_{i}\widehat{p}_{i}^{{}^{\prime}}\left(y_{i}-\widehat{G}\left(\widehat{s}_{i}\right)\right)^{2},\ \widehat{\Sigma}_{1}\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{\overline{\mu}}_{i}^{I}\widehat{\overline{\mu}}_{i}^{I^{\prime}},$
$\widehat{\overline{\mu}}_{i}^{I}$ $\coloneqq n^{-1}\sum_{j=1}^{n}\widehat{G}_{V}\left(\widehat{s}_{j}\right)\widehat{p}_{j}q_{j}^{{}^{\prime}}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{ and }\widehat{v}_{ji}\coloneqq\left(\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-\widehat{F}\left(\left.x_{j}\right|z_{i},w_{i}\right)\right).$
IN02 have proved asymptotic normality for known and scalar-valued functionals of $G\left(s\right)$. I apply their results to $c^{{}^{\prime}}\widehat{b}_{1}\left(v,w\right)$ for any constant vector $c^{{}^{\prime}}c=1$ and obtain
@@
By (A.15) and Assumption 7(f),
[…]
which implies
[…]
Then,
$\ n^{1/2}\widehat{\Omega}_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
$=$ $\ \left(c_{1n}\widehat{\Omega}_{1}\right)^{-1/2}\left(c_{1n}\Omega_{1}\right)^{1/2}n^{1/2}\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
[…]
Proof of Theorem 3.
Following the proof of Lemma 2, under regularity conditions one may use the Cramér–Wold device to extend the results from scalar-valued functionals to vector-valued functionals of $G\left(s\right)$. Therefore, I maintain that $b\left(x\right)$ is a scalar in this proof. First, I derive the influence functions for $\widehat{b}\left(x\right)$ that correctly account for the estimation errors from each step and prove its asymptotic normality. Then, I show consistency for the estimator of the variance of $\widehat{b}\left(x\right)$, which can be used to construct asymptotically valid confidence intervals. I write $r^{M_{3}}\left(x\right)$ as $r\left(x\right)$ when there is no confusion. The proof relies on the results from IN02, and I point out differences.
Define
$\Omega_{2}$ $\coloneqq\Omega_{21}+\Omega_{22},\ \Omega_{21}\coloneqq\mathbb{E}\left(A_{1}P^{-1}p_{i}u_{i}\right)\left(A_{1}P^{-1}p_{i}u_{i}\right)^{{}^{\prime}},$ (A.19)
$\Omega_{22}$ $\coloneqq\mathbb{E}\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r^{M_{3}}\left(x\right)^{{}^{\prime}}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)^{{}^{\prime}},$
$A_{1}$ $\coloneqq r^{M_{3}}\left(x\right)^{{}^{\prime}}\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}},\ \overline{\mu}_{i}^{II}\coloneqq\mathbb{E}\left[\left.r_{j}b_{1,V}\left(V_{j},W_{j}\right)q_{j}^{{}^{\prime}}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\text{ and}$
$b_{1,V}\left(V_{j},W_{j}\right)$ $\coloneqq\left.\partial b_{1}\left(v,w\right)/\partial v\right|_{v=V_{j},w=W_{j}},$
and $F\coloneqq\Omega_{2}^{-1/2}$, which is well-defined because $\Omega_{2}=\Omega_{21}+\Omega_{22}$ and
$\Omega_{21}$ $=A_{1}P^{-1}\left(\mathbb{E}p_{i}p_{i}^{{}^{\prime}}u_{i}^{2}\right)P^{-1}A_{1}^{{}^{\prime}}$
$=A_{1}P^{-1}\left(\mathbb{E}\left[p_{i}p_{i}^{{}^{\prime}}\mathbb{E}\left(\left.u_{i}^{2}\right|X_{i},Z_{i},W_{i}\right)\right]\right)P^{-1}A_{1}^{{}^{\prime}}$
$\geq CA_{1}P^{-1}A_{1}^{{}^{\prime}}=Cr\left(x\right)^{{}^{\prime}}\left(\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right)P^{-1}\left(\mathbb{E}\overline{p}_{i}r_{i}^{{}^{\prime}}\right)r\left(x\right)>0,$ (A.20)
Define functionals:
$\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)$ $\coloneqq\widehat{\mathbb{E}}\left[\left.\widehat{b}_{1}\left(\widehat{V},W\right)\right|X=x\right]=\widehat{b}\left(x\right),$
$\widehat{a}\left(b_{1},\widehat{V}\right)$ $\coloneqq\widehat{\mathbb{E}}\left[\left.b_{1}\left(\widehat{V},W\right)\right|X=x\right],$
@@
$\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$
$=$ $\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},\widehat{V}\right)+\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)+\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
@@
Let $\widehat{H}_{1}\coloneqq F\widehat{A}_{1}\widehat{P}^{-1}$, $\widehat{A}_{1}\coloneqq r\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}n^{-1}\sum_{i=1}^{n}r_{i}\widehat{\overline{p}}_{i}$, $\widehat{R}\coloneqq n^{-1}\sum_{i=1}^{n}r_{i}r_{i}^{{}^{\prime}}$, $\widehat{\overline{p}}_{i}\coloneqq\overline{p}\left(\widehat{s}_{i}\right)$, $\widehat{H}_{2}\coloneqq FA_{2}\widehat{R}^{-1}$, $G\coloneqq\left(G\left(s_{1}\right),...,G\left(s_{n}\right)\right)^{{}^{\prime}}$, and $\widetilde{G}\coloneqq\left(G\left(\widehat{s}_{1}\right),...,G\left(\widehat{s}_{n}\right)\right)^{{}^{\prime}}$. First, for $\psi_{1i}$,
$\ n^{1/2}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},\widehat{V}\right)\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(\widehat{B}-\widetilde{B}\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(n^{-1}\widehat{\overline{p}}\widehat{P}^{-1}\widehat{p}^{{}^{\prime}}Y-\widetilde{B}\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left[n^{-1}\widehat{\overline{p}}\widehat{P}^{-1}\widehat{p}^{{}^{\prime}}\left(Y-G+G-\widetilde{G}+\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)+\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right]$
$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{1}\widehat{p}_{i}\left[u_{i}-\left(G\left(\widehat{s}_{i}\right)-G\left(s_{i}\right)\right)\right]+n^{-1/2}\widehat{H}_{1}\widehat{p}^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)$
$\ +n^{-1/2}\widehat{H}_{2}r^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)$ (A.23)
@@
because $\left\|H_{1}\right\|^{2}\leq CA_{1}A_{1}^{{}^{\prime}}/\Omega_{2}\leq C$ and $\left\|H_{2}\right\|^{2}=A_{2}A_{2}^{{}^{\prime}}/\Omega_{2}\leq CA_{1}A_{1}^{{}^{\prime}}/\Omega_{2}\leq C$. In addition, $\left\|\widehat{P}-P\right\|=o_{p}\left(1\right)$, $\left\|\widehat{R}-I\right\|=o_{p}\left(1\right)$, and $\left\|n^{-1}\sum_{i}r_{i}\widehat{\overline{p}}_{i}-\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right\|=o_{p}\left(1\right)$ by Lemma A3 of IN02 and Newey (1997). I also know that $\lambda_{\min}\left(P\right)\geq C>0$ and $\left\|P\right\|=O\left(1\right)$ by Assumption 5. By Slutsky’s theorem, $\left\|\widehat{R}^{-1}-I\right\|=o_{p}\left(1\right)$ and $\left\|\widehat{P}^{-1}-P^{-1}\right\|=o_{p}\left(1\right)$. By the CS inequality and Lemma A3 of IN02,
$\left\|n^{-1}\sum_{i}r_{i}\left(\overline{p}_{i}-\widehat{\overline{p}}_{i}\right)^{{}^{\prime}}\right\|^{2}$ $\leq n^{-1}\sum_{i}\left\|r_{i}\right\|^{2}\times n^{-1}\sum_{i}\left\|\widehat{\overline{p}}_{i}-\overline{p}_{i}\right\|^{2}$
@@
$=$ $\ 2\left\|F\left(r\left(x\right)^{{}^{\prime}}\left(I+o_{p}\left(1\right)\right)\left(\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}+o_{p}\left(1\right)\right)-r\left(x\right)^{{}^{\prime}}\mathbb{E}r_{i}\overline{p}_{i}^{{}^{\prime}}\right)\widehat{P}^{-1}\right\|^{2}$
@@
$n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)=$ $\ O_{p}\left(M_{2}^{-2d_{4}/j_{4}}\right),\text{ and}$
$n^{-1}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)=$ $\ O_{p}\left(M_{2}^{-2d_{4}/j_{4}}\right).$
Therefore, for $D_{12}$,
$\left|n^{-1/2}\widehat{H}_{1}\widehat{p}^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right|^{2}$ $\leq n\left[\widehat{H}_{1}\widehat{P}\widehat{H}_{1}^{{}^{\prime}}\right]\left[n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{{}^{\prime}}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right]$
@@
For $D_{13}$, similarly to (A.29),
$\left|n^{-1/2}\widehat{H}_{2}r^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right|^{2}$ $\leq n\left[\widehat{H}_{2}\widehat{R}\widehat{H}_{2}\right]\left[n^{-1}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)^{{}^{\prime}}\left(\widehat{\overline{p}}\overline{\alpha}^{M_{2}}-\widetilde{B}\right)\right]$
@@
To prove $\psi_{2i}=H_{2}\overline{\mu}_{i}^{II}$, first notice that by Taylor expansion,
$\ n^{1/2}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{{}^{\prime}}\widehat{R}^{-1}r^{{}^{\prime}}\left(\widetilde{B}-B\right)$
@@
$\delta_{ij}=F_{\left.X_{i}\right|Z_{j},W_{j}}\left(\left.x_{i}\right|z_{j},w_{j}\right)-q_{j}^{{}^{\prime}}\gamma^{M_{1}}\left(x_{i}\right),\ \Delta_{i}^{I}=q_{i}^{{}^{\prime}}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}v_{ij},$
$\Delta_{i}^{II}=q_{i}^{{}^{\prime}}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}\delta_{ij},\text{ and }\Delta_{i}^{III}=-\delta_{ii}.$ (A.34)
Following an argument identical to the proof of Lemma B7 of IN02,
@@
$\left|D_{212}\right|^{2}$ $\leq Cn\left[\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{{}^{\prime}}\right]\left[n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right]$
@@
$\left|D_{213}\right|^{2}\leq Cn\left[H_{2}\widehat{R}H_{2}^{{}^{\prime}}\right]\left[n^{-1}\sum_{i}\left(\left(\Delta_{i}^{II}\right)^{2}+\left(\Delta_{i}^{III}\right)^{2}\right)\right]=O_{p}\left(nM_{1}^{1-2d_{1}/j_{1}}\right)=o_{p}\left(1\right),$ (A.37)
where the first equality is established in the proof of Theorem 4 of IN02.
@@
$\left|D_{22}\right|$ $\leq Cn^{1/2}\left\|\widehat{H}_{2}\right\|\sup_{x\in\mathcal{X}}\left\|r\left(x\right)\right\|\left|n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right|$
$=O_{p}\left(n^{1/2}\zeta\left(M_{3}\right)\Delta_{1n}^{2}\right)=o_{p}\left(1\right).$ (A.38)
@@
$n^{1/2}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}\overline{\mu}_{i}^{II}+o_{p}\left(1\right).$ (A.39)
@@
$\ n^{1/2}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}b_{1}\left(v_{i},w_{i}\right)-n^{1/2}Fb\left(x\right)$
@@
$+n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}\left(b\left(x_{i}\right)-r_{i}^{{}^{\prime}}\rho^{M_{3}}\right)-n^{1/2}F\left(b\left(x\right)-r\left(x\right)^{{}^{\prime}}\rho^{M_{3}}\right)$
$\eqqcolon$ $\ D_{31}+D_{32}+D_{33}+D_{34},$

### v3

@@ A.1 Proofs
$\overline{\mu}_{i}^{I}$ $\coloneqq\mathbb{E}\left[\left.G_{V}\left(S_{j}\right)\tau_{V}\left(V_{j}\right)p_{j}q_{j}^{\prime}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\ u_{i}\coloneqq Y_{i}-G\left(S_{i}\right),$
@@
$\widehat{\Omega}_{1}$ $\coloneqq\overline{p}^{M_{2}}\left(v,w\right)^{\prime}\widehat{P}^{-1}\left(\widehat{\Sigma}+\widehat{\Sigma}_{1}\right)\widehat{P}^{-1}\overline{p}^{M_{2}}\left(v,w\right),$ (A.13)
$\widehat{\Sigma}$ $\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{p}_{i}\widehat{p}_{i}^{\prime}\left(y_{i}-\widehat{G}\left(\widehat{s}_{i}\right)\right)^{2},\ \widehat{\Sigma}_{1}\coloneqq n^{-1}\sum_{i=1}^{n}\widehat{\overline{\mu}}_{i}^{I}\widehat{\overline{\mu}}_{i}^{I}{}^{\prime},$
$\widehat{\overline{\mu}}_{i}^{I}$ $\coloneqq n^{-1}\sum_{j=1}^{n}\widehat{G}_{V}\left(\widehat{s}_{j}\right)\widehat{p}_{j}q_{j}^{\prime}\widehat{Q}^{-1}q_{i}\widehat{v}_{ji},\text{ and }\widehat{v}_{ji}\coloneqq\left(\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-\widehat{F}\left(\left.x_{j}\right|z_{i},w_{i}\right)\right).$
IN02 have proved asymptotic normality for known and scalar-valued functionals of $G\left(s\right)$. I apply their results to $c^{\prime}\widehat{b}_{1}\left(v,w\right)$ for any constant vector $c^{\prime}c=1$ and obtain
$c^{\prime}\sqrt{n}\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$ $\overset{d}{\longrightarrow}N\left(0,1\right)\ \text{and}$
$\left(c^{\prime}\Omega_{1}c\right)^{-1}\left[c^{\prime}\left(\widehat{\Omega}_{1}-\Omega_{1}\right)c\right]$ $\overset{p}{\longrightarrow}0.$ (A.14)
By (A.14) and Assumption 7(f),
$c^{\prime}\left(c_{1n}\widehat{\Omega}_{1}-c_{1n}\Omega_{1}\right)c\overset{p}{\longrightarrow}0,$ (A.15)
which implies
$c_{1n}\widehat{\Omega}_{1}\overset{p}{\longrightarrow}\overline{\Omega}_{1}.$ (A.16)
Then,
$\ \sqrt{n}\widehat{\Omega}_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
$=$ $\ \left(c_{1n}\widehat{\Omega}_{1}\right)^{-1/2}\left(c_{1n}\Omega_{1}\right)^{1/2}\sqrt{n}\Omega_{1}^{-1/2}\left(\widehat{b}_{1}\left(v,w\right)-b_{1}\left(v,w\right)\right)$
$\overset{d}{\longrightarrow}$ $\ \overline{\Omega}_{1}^{-1/2}\overline{\Omega}_{1}^{1/2}N\left(0,I\right)=_{d}N\left(0,I\right),$ (A.17)
Proof of Theorem 3.
Similarly to the proof of Lemma 2, by the Cramér–Wold device it suffices to consider the case when $b\left(x\right)$ is a scalar. First, I derive the influence functions for $\widehat{b}\left(x\right)$ that correctly account for the estimation errors from each step and prove its asymptotic normality. Then, I show consistency for the estimator of the variance of $\widehat{b}\left(x\right)$. I write $r^{M_{3}}\left(x\right)$ as $r\left(x\right)$ when there is no confusion. The proof relies on the results from IN02, and I point out differences.
Define
$\Omega_{2}$ $\coloneqq\Omega_{21}+\Omega_{22},\text{ where}$
$\Omega_{21}$ $\coloneqq\mathbb{E}\left(A_{1}P^{-1}p_{i}u_{i}\right)\left(A_{1}P^{-1}p_{i}u_{i}\right)^{\prime},$
$\Omega_{22}$ $\coloneqq\mathbb{E}\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r\left(x\right)^{\prime}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)\left(A_{1}P^{-1}\overline{\mu}_{i}^{I}-r\left(x\right)^{\prime}\left(\overline{\mu}_{i}^{II}+r_{i}\xi_{i}\right)\right)^{\prime},$
$A_{1}$ $\coloneqq r\left(x\right)^{\prime}\mathbb{E}r_{i}\overline{p}_{i}^{\prime},$
$\overline{\mu}_{i}^{II}$ $\coloneqq\mathbb{E}\left[\left.r_{j}b_{1,V}\left(V_{j},W_{j}\right)q_{j}^{\prime}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],\text{ and}$
$b_{1,V}\left(V_{j},W_{j}\right)$ $\coloneqq\left.\partial b_{1}\left(v,w\right)/\partial v\right|_{v=V_{j},w=W_{j}}.$ (A.18)
Let $F\coloneqq\Omega_{2}^{-1/2}$, which is well-defined because $\Omega_{2}=\Omega_{21}+\Omega_{22}$ and
$\Omega_{21}$ $=A_{1}P^{-1}\left(\mathbb{E}p_{i}p_{i}^{\prime}u_{i}^{2}\right)P^{-1}A_{1}^{\prime}$
$=A_{1}P^{-1}\left(\mathbb{E}\left[p_{i}p_{i}^{\prime}\mathbb{E}\left(\left.u_{i}^{2}\right|X_{i},Z_{i},W_{i}\right)\right]\right)P^{-1}A_{1}^{\prime}$
$\geq CA_{1}P^{-1}A_{1}^{\prime}=Cr\left(x\right)^{\prime}\left(\mathbb{E}r_{i}\overline{p}_{i}^{\prime}\right)P^{-1}\left(\mathbb{E}\overline{p}_{i}r_{i}^{\prime}\right)r\left(x\right)>0,$ (A.19)
Define functionals:
@@
$\ \sqrt{n}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-a\left(b_{1},V\right)\right)$
$=$ $\ \sqrt{n}F\left(\widehat{a}\left(\widehat{b}_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},\widehat{V}\right)+\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)+\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
@@
where $H_{1}\coloneqq FA_{1}P^{-1}$, $A_{1}=r\left(x\right)^{\prime}R^{-1}\mathbb{E}r_{i}\overline{p}_{i}^{\prime}=r\left(x\right)^{\prime}\mathbb{E}r_{i}\overline{p}_{i}^{\prime}$, $r_{i}\coloneqq r\left(x_{i}\right)$, $P\coloneqq\mathbb{E}p_{i}p_{i}^{\prime}$, $p_{i}\coloneqq p^{M_{2}}\left(X_{i},V_{i},W_{i}\right)$, $u_{i}\coloneqq y_{i}-G\left(s_{i}\right)$, $\text{$\overline{\mu}_{i}^{I}$}\coloneqq\mathbb{E}\left[\left.G_{V}\left(S_{j}\right)\tau_{V}\left(V_{j}\right)p_{j}q_{j}^{\prime}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right],$ $G_{V}\left(S_{j}\right)\coloneqq\left.\partial G\left(s\right)/\partial v\right|_{s=S_{j}},$ $v_{ji}\coloneqq\mathbf{\mathbbm{1}}\left\{x_{i}\leq x_{j}\right\}-F\left(\left.x_{j}\right|z_{i},w_{i}\right)$, $H_{2}\coloneqq FA_{2}R^{-1}=FA_{2}$, $A_{2}\coloneqq r\left(x\right)^{\prime}$, $\overline{\mu}_{i}^{II}\coloneqq\mathbb{E}\left[\left.r_{j}b_{1,V}\left(V_{j},W_{j}\right)q_{j}^{\prime}Q^{-1}q_{i}v_{ji}\right|\mathcal{I}_{i}\right]$, $b_{1,V}\left(V_{j},W_{j}\right)=\left.\partial b_{1}\left(v,w\right)/\partial v\right|_{v=V_{j}},$ and $\xi_{i}=b_{1}\left(v_{i},w_{i}\right)-b\left(x_{i}\right)$.
Let $\widehat{H}_{1}\coloneqq F\widehat{A}_{1}\widehat{P}^{-1}$, $\widehat{A}_{1}\coloneqq r\left(x\right)^{\prime}\widehat{R}^{-1}n^{-1}\sum_{i=1}^{n}r_{i}\widehat{\overline{p}}_{i}^{\prime}$, $\widehat{R}\coloneqq n^{-1}\sum_{i=1}^{n}r_{i}r_{i}^{\prime}$, $\widehat{\overline{p}}_{i}\coloneqq\overline{p}\left(\widehat{s}_{i}\right)$, $\widehat{H}_{2}\coloneqq FA_{2}\widehat{R}^{-1}$, $G\coloneqq\left(G\left(s_{1}\right),\ldots,G\left(s_{n}\right)\right)^{\prime}$, and $\widetilde{G}\coloneqq\left(G\left(\widehat{s}_{1}\right),\ldots,G\left(\widehat{s}_{n}\right)\right)^{\prime}$. First, for $\psi_{1i}$,
@@
[…]
$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{1}\widehat{p}_{i}\left[u_{i}-\left(G\left(\widehat{s}_{i}\right)-G\left(s_{i}\right)\right)\right]+n^{-1/2}\widehat{H}_{1}\widehat{p}^{\prime}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)$
[…]
@@
because $\left\|H_{1}\right\|^{2}\leq CA_{1}A_{1}^{\prime}/\Omega_{2}\leq C$ and $\left\|H_{2}\right\|^{2}=A_{2}A_{2}^{\prime}/\Omega_{2}\leq CA_{1}A_{1}^{\prime}/\Omega_{2}\leq C$. In addition, $\left\|\widehat{P}-P\right\|=o_{p}\left(1\right)$, $\left\|\widehat{R}-I\right\|=o_{p}\left(1\right)$, and $\left\|n^{-1}\sum_{i}r_{i}\widehat{\overline{p}}_{i}^{\prime}-\mathbb{E}r_{i}\overline{p}_{i}^{\prime}\right\|=o_{p}\left(1\right)$ by Lemma A3 of IN02 and Newey (1997). Furthermore, $\lambda_{\min}\left(P\right)\geq C>0$ and $\left\|P\right\|=O\left(1\right)$ by Assumption 5. By Slutsky’s theorem, $\left\|\widehat{R}^{-1}-I\right\|=o_{p}\left(1\right)$ and $\left\|\widehat{P}^{-1}-P^{-1}\right\|=o_{p}\left(1\right)$. By the CS inequality and Lemma A3 of IN02,
$\left\|n^{-1}\sum_{i}r_{i}\left(\overline{p}_{i}-\widehat{\overline{p}}_{i}\right)^{\prime}\right\|^{2}$ $\leq n^{-1}\sum_{i}\left\|r_{i}\right\|^{2}\times n^{-1}\sum_{i}\left\|\widehat{\overline{p}}_{i}-\overline{p}_{i}\right\|^{2}$
@@
$=$ $\ 2\left\|F\left(r\left(x\right)^{\prime}\left(I+o_{p}\left(1\right)\right)\left(\mathbb{E}r_{i}\overline{p}_{i}^{\prime}+o_{p}\left(1\right)\right)-r\left(x\right)^{\prime}\mathbb{E}r_{i}\overline{p}_{i}^{\prime}\right)\widehat{P}^{-1}\right\|^{2}$
@@
$n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{\prime}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)=$ $\ O_{p}\left(M_{2}^{-2d_{4}/j_{4}}\right)\text{ and}$
[…]
Therefore, for $D_{12}$, I have
$\left|n^{-1/2}\widehat{H}_{1}\widehat{p}^{\prime}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right|^{2}$ $\leq n\left[\widehat{H}_{1}\widehat{P}\widehat{H}_{1}^{\prime}\right]\left[n^{-1}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)^{\prime}\left(\widetilde{G}-\widehat{p}\overline{\alpha}^{M_{2}}\right)\right]$
@@
For $D_{13}$, similarly to (A.27), I have
[…]
@@
To prove $\psi_{2i}=H_{2}\overline{\mu}_{i}^{II}$, by Taylor expansion it is true that
$\ \sqrt{n}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)$
$=$ $\ n^{-1/2}Fr\left(x\right)^{\prime}\widehat{R}^{-1}r^{\prime}\left(\widetilde{B}-B\right)$
@@
$\delta_{ij}=F_{\left.X_{i}\right|Z_{j},W_{j}}\left(\left.x_{i}\right|z_{j},w_{j}\right)-q_{j}^{\prime}\gamma^{M_{1}}\left(x_{i}\right),\ \Delta_{i}^{I}=q_{i}^{\prime}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}v_{ij},$
$\Delta_{i}^{II}=q_{i}^{\prime}\widehat{Q}^{-1}n^{-1}\sum_{j}q_{j}\delta_{ij},\text{ and }\Delta_{i}^{III}=-\delta_{ii}.$ (A.32)
Following an argument identical to the proof of Lemma B7 in IN02,
@@
$\left|D_{212}\right|^{2}$ $\leq Cn\left[\left(\widehat{H}_{2}-H_{2}\right)\widehat{R}\left(\widehat{H}_{2}-H_{2}\right)^{\prime}\right]\left[n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right]$
@@
$\left|D_{213}\right|^{2}\leq Cn\left[H_{2}\widehat{R}H_{2}^{\prime}\right]\left[n^{-1}\sum_{i}\left(\left(\Delta_{i}^{II}\right)^{2}+\left(\Delta_{i}^{III}\right)^{2}\right)\right]=O_{p}\left(nM_{1}^{1-2d_{1}/j_{1}}\right)=o_{p}\left(1\right),$ (A.35)
where the first equality is established in the proof of Theorem 4 in IN02.
@@
$\left|D_{22}\right|$ $\leq C\sqrt{n}\left\|\widehat{H}_{2}\right\|\sup_{x\in\mathcal{X}}\left\|r\left(x\right)\right\|\left|n^{-1}\sum_{i}\left(\widehat{v}_{i}-v_{i}\right)^{2}\right|$
$=O_{p}\left(\sqrt{n}\zeta\left(M_{3}\right)\Delta_{1n}^{2}\right)=o_{p}\left(1\right).$ (A.36)
@@
$\sqrt{n}F\left(\widehat{a}\left(b_{1},\widehat{V}\right)-\widehat{a}\left(b_{1},V\right)\right)=n^{-1/2}\sum_{i}H_{2}\overline{\mu}_{i}^{II}+o_{p}\left(1\right).$ (A.37)
@@
$\ \sqrt{n}F\left(\widehat{a}\left(b_{1},V\right)-a\left(b_{1},V\right)\right)$
$=$ $\ n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}b_{1}\left(v_{i},w_{i}\right)-\sqrt{n}Fb\left(x\right)$
@@
$+n^{-1/2}\sum_{i}\widehat{H}_{2}r_{i}\left(b\left(x_{i}\right)-r_{i}^{\prime}\rho^{M_{3}}\right)-\sqrt{n}F\left(b\left(x\right)-r\left(x\right)^{\prime}\rho^{M_{3}}\right)$
$\eqqcolon$ $\ D_{31}+D_{32}+D_{33}+D_{34},$ (A.38)